OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. a. y = -x + 6 b. y = x - 6 c. y = -x - 2 d. y = x + 2

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

thank youuu

OpenStudy (anonymous):

i can teach you a method on how to check perpendicular means the slope is negative reciprocal. since orig equation has -1 as slope the perpendicular line has 1 as slope then it passes through (-4,-2) just sub into the equation y = x + 2 -2 = -4 + 2 =2 = =2 therefore, it can be instantaneously checked